A Complete Bibliography of Publications in Communications of the AMS: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

13 August 2025 Version 1.00

Title word cross-reference

B + B + t [33]. C^{0} [37]. K(1) [14]. L^{2}(Γ \ G) [23]. L^{∞} [9]. m = 2 [19]. ∫↕_{N} [15]. p [36].

-equivariant [3]. -local [14].

1D [30].

4-manifolds [3].

algebraic [4]. algebraically [36]. algebras [13]. alternating [26]. Ampère [37]. amplituhedron [19]. analysis [38, 2]. Anderson [43]. Anosov [38]. Antiholomorphic [34]. approach [27]. arbitrary [29]. assuming [26]. Asymptotic [4]. Automorphic [24]. automorphisms [28, 18].

Bäcklund [7]. band [38]. based [40]. beyond [39]. bipartite [12]. bisections [11]. bootstrap [24]. boundary [35]. braid [25].

Calogero [47]. Catalan [31]. centroaffine [7]. chaos [32]. closed [36]. clusters [19]. codimension [30]. coefficients [48]. Cohen [45]. cohomology [36]. Colored [20]. concordance [10]. conditions [9]. conformal [24]. conjecture [33, 26]. connectivity [35]. contact [38]. continuation [48]. continuum [47]. convergence [26]. convex [37, 40]. convexity [37, 27]. Conze [29]. correspondences [34]. critical [47]. cubic [1, 30]. curve [35]. curves [7, 22]. descent [40]. deterministic [39]. dichotomy [41]. Different [28]. dimension [42]. dimensional [2]. domains [37]. duality [9]. dynamical [32, 8].

eigenfunctions [23]. Eigenvalue [9, 2]. Eisenstein [49]. energy [42]. enforcing [37]. enveloping [13]. equation [7, 37, 30]. equations [13]. equivariant [3]. Erdos [33, 26]. Ergodic [41]. error [8, 2]. Eulerian [19]. existence [45]. expansions [42]. Expressive [22].

fast [39, 5]. fast-slow [39]. fermionic [20]. fields [36]. First [44, 17]. first-passage [17]. flows [38, 41]. focusing [30]. forms [3]. framework [8]. Friendly [11]. fronts [6]. functions [20].

Gap [1, 46]. generalization [2]. Geodesic [17]. geometric [27]. geometry [7]. Gevrey [48]. Gordon [30]. gradient [40]. graph [35]. graphs [12, 28, 11, 1]. Gromov [35]. Grothendieck [14]. group [13]. groups [44]. growth [16].

Hardy [26]. high [2]. Higher [14, 23, 41]. Hochschild [36]. homeomorphism [44]. homology [36]. homomorphisms [28]. hyperbolic [18]. hypersimplex [19].

II [15]. interior [37]. Intersection [3]. invariant [3].

Kazhdan [25]. Klein [30]. knots [10].

Lamé [7]. last [32]. layer [2]. learning [8]. least [37]. least-squares [37]. length [17]. Lesigne [29]. limits [39]. line [47]. linear [46, 13, 35]. Littlewood [26]. local [14, 15, 38, 45, 18]. localization [43]. Low [42]. Lusztig [25].

Macaulay [45]. machine [8]. Mahowald [3]. manifolds [3, 44]. Marcus [39]. mating [34]. matroids [25]. Maximum [12]. method [37, 40]. Micro [38]. Micro-local [38]. mod [36]. mod- [36]. model [8]. models [20, 47]. modules [15, 45]. Monge [37]. Moser [47]. motivic [36].

networks [2]. neural [2]. Non [43, 45, 40]. non-convex [40]. Non-existence [45]. Non-stationary [43]. nonlinear [37]. numbers [31, 19]. nutrients [16].

one [30]. operators [48]. optimal [9]. optimality [9]. optimization [40]. order [44]. ordering [10].

partial [10]. partially [48]. passage [32, 17]. penalty [37]. percolation [32, 17]. perspectives [28]. perversity [21]. Pin [3]. planar [37]. polynomials [25]. posedness [47]. positive [23]. Positroid [31]. presence [27]. prime [26]. priori [2]. problems [9, 2]. processes [46]. projective [46]. proof [33]. pulled [6].

Quantum [28, 15]. quasi [13]. quasi-linear [13].

random [11]. rank [23, 41]. ray [27]. Realization [34]. regime [39]. Regularity [16]. relations [9]. Relative [21]. resolvent [42]. Ribbon [10]. rigidity [18, 44]. rings [45]. Ruelle [38].

Scaling [47]. Scaling-critical [47]. scenario [5]. Schrödinger [48, 2]. selection [6]. Self [7]. Self-Bäcklund [7]. semiadditive [14]. semiclassical [27]. series [26]. sets [1]. shifted [17]. Signs [19]. slow [39]. Smooth [18, 37]. soliton [30]. solving [2]. SPDEs [46]. spectra [1, 24]. Spectral [46]. spectrum [38, 49]. sphere [14]. Spheres [35]. spin [3]. spread [12]. spreading [5]. squares [37]. Stability [32, 16, 30]. stationary [43]. strictly [37]. structure [38, 29, 13]. subgroups [15]. subspace [41]. Superdiffusive [39]. Swarm [40]. Swarm-based [40]. symmetric [20]. Symmetries [15]. systems [39, 29, 8]. syzygies [4].

Temperedness [23]. theorems [34]. theory [14]. tilings [19]. toral [18]. transforms [27]. transport [9]. Tumor [16]. tuples [26]. two [42, 2]. two-layer [2]. Type [15].

Ulrich [45]. Unique [48]. Universal [6, 13]. unramified [49].

vanishing [4]. varieties [4]. vertex [20]. via [13].

weights [17]. well [47]. well-posedness [47]. Witt [14].

X [27]. X-ray [27].

References

Kollar:2021:GSS

[1] Alicia Kollár and Peter Sarnak. Gap sets for the spectra of cubic graphs. Communications of the AMS, 1(1):1–38, October 2021. ISSN 2692-3688.

Lu:2022:PGE

[2] Jianfeng Lu and Yulong Lu. A priori generalization error analysis of two-layer neural networks for solving high dimensional Schrödinger eigenvalue problems. Communications of the AMS, 2(1):1–21, January 2022. ISSN 2692-3688.

Hopkins:2022:IFS

[3] Michael Hopkins, Jianfeng Lin, XiaoLin Danny Shi, and Zhouli Xu. Intersection forms of spin 4-manifolds and the Pin(2)-equivariant Mahowald invariant. Communications of the AMS, 2(2): 22–132, February 2022. ISSN 2692-3688.

Park:2022:AVS

[4] Jinhyung Park. Asymptotic vanishing of syzygies of algebraic varieties. Communications of the AMS, 2(3):133–148, June 2022. ISSN 2692-3688.

He:2022:FSS

[5] Siming He, Eitan Tadmor, and Andrej Zlato. On the fast spreading scenario. Communications of the AMS, 2(4):149– 171, July 2022. ISSN 2692-3688.

Avery:2022:USP

[6] Montie Avery and Arnd Scheel. Universal selection of pulled fronts. Communications of the AMS, 2(5):172–231, July 2022. ISSN 2692-3688.

Bialy:2022:SBC

[7] Misha Bialy, Gil Bor, and Serge Tabachnikov. Self-Bäcklund curves in centroaffine geometry and Lamé’s equation. Communications of the AMS, 2(6):232– 282, August 2022. ISSN 2692-3688.

Levine:2022:FML

[8] Matthew Levine and Andrew Stuart. A framework for machine learning of model error in dynamical systems. Communications of the AMS, 2(7):283–344, October 2022. ISSN 2692-3688.

Bungert:2022:EPO

[9] Leon Bungert and Yury Korolev. Eigenvalue problems in L^{∞}: optimality conditions, duality, and relations with optimal transport. Communications of the AMS, 2(8):345–373, October 2022. ISSN 2692-3688.

Agol:2022:RCK

[10] Ian Agol. Ribbon concordance of knots is a partial ordering. Communications of the AMS, 2(9):374–379, December 2022. ISSN 2692-3688.

Ferber:2022:FBR

[11] Asaf Ferber, Matthew Kwan, Bhargav Narayanan, Ashwin Sah, and Mehtaab Sawhney. Friendly bisections of random graphs. Communications of the AMS, 2 (10):380–416, December 2022. ISSN 2692-3688.

Breen:2022:MSG

[12] Jane Breen, Alex W. N. Riasanovsky, Michael Tait, and John Urschel. Maximum spread of graphs and bipartite graphs. Communications of the AMS, 2 (11):417–480, December 2022. ISSN 2692-3688.

Linares:2023:SGQ

[13] Pablo Linares, Felix Otto, and Markus Tempelmayr. The structure group for quasi-linear equations via universal enveloping algebras. Communications of the AMS, 3(1):1–64, January 2023. ISSN 2692-3688.

Carmeli:2023:HSG

[14] Shachar Carmeli and Allen Yuan. Higher semiadditive Grothendieck–Witt theory and the K(1)-local sphere. Communications of the AMS, 3(2):65–111, March 2023. ISSN 2692-3688.

Edie-Michell:2023:TIQ

[15] Cain Edie-Michell. Type II quantum subgroups of ∫↕_{N}. I: Symmetries of local modules. Communications of the AMS, 3 (3):112–165, May 2023. ISSN 2692-3688.

Jacobs:2023:TGN

[16] Matt Jacobs, Inwon Kim, and Jiajun Tong. Tumor growth with nutrients: Regularity and stability. Communications of the AMS, 3(4):166–208, May 2023. ISSN 2692-3688.

Krishnan:2023:GLS

[17] Arjun Krishnan, Firas Rassoul-Agha, and Timo Seppäläinen. Geodesic length and shifted weights in first-passage percolation. Communications of the AMS, 3(5): 209–289, May 2023. ISSN 2692-3688.

Kalinin:2023:SLR

[18] Boris Kalinin, Victoria Sadovskaya, and Zhenqi Wang. Smooth local rigidity for hyperbolic toral automorphisms. Communications of the AMS, 3(6):290–328, June 2023. ISSN 2692-3688.

Parisi:2023:AHS

[19] Matteo Parisi, Melissa Sherman-Bennett, and Lauren Williams. The m = 2 amplituhedron and the hypersimplex: Signs, clusters, tilings, Eulerian numbers. Communications of the AMS, 3(7):329–399, July 2023. ISSN 2692-3688.

Aggarwal:2023:CFV

[20] Amol Aggarwal, Alexei Borodin, and Michael Wheeler. Colored fermionic vertex models and symmetric functions. Communications of the AMS, 3(08):400– 630, August 2023. ISSN 2692-3688.

Hansen:2023:RP

[21] David Hansen and Peter Scholze. Relative perversity. Communications of the AMS, 3(9):631–668, August 2023. ISSN 2692-3688.

Fomin:2023:EC

[22] Sergey Fomin and Eugenii Shustin. Expressive curves. Communications of the AMS, 3(10):669–743, August 2023. ISSN 2692-3688.

Edwards:2023:TPE

[23] Sam Edwards and Hee Oh. Temperedness of L^{2}(Γ \ G) and positive eigenfunctions in higher rank. Communications of the AMS, 3(11):744–778, September 2023. ISSN 2692-3688.

Kravchuk:2024:ASC

[24] Petr Kravchuk, Dalimil Mazá, and Sridip Pal. Automorphic spectra and the conformal bootstrap. Communications of the AMS, 4(1):1–63, January 2024. ISSN 2692-3688.

Ferroni:2024:KLP

[25] Luis Ferroni and Matt Larson. Kazhdan– Lusztig polynomials of braid matroids. Communications of the AMS, 4(2):64–79, January 2024. ISSN 2692-3688.

Tao:2024:CAS

[26] Terence Tao. The convergence of an alternating series of Erdős, assuming the Hardy–Littlewood prime tuples conjecture. Communications of the AMS, 4(03): 80–96, February 2024. ISSN 2692-3688.

Vasy:2024:SAG

[27] András Vasy. A semiclassical approach to geometric X-ray transforms in the presence of convexity. Communications of the AMS, 4(4):97–116, February 2024. ISSN 2692-3688.

Daws:2024:QGD

[28] Matthew Daws. Quantum graphs: Different perspectives, homomorphisms and quantum automorphisms. Communications of the AMS, 4(5):117–181, February 2024. ISSN 2692-3688.

Jamneshan:2024:SAC

[29] Asgar Jamneshan, Or Shalom, and Terence Tao. The structure of arbitrary Conze–Lesigne systems. Communications of the AMS, 4(6):182–229, February 2024. ISSN 2692-3688.

Luhrmann:2024:COS

[30] Jonas Lührmann and Wilhelm Schlag. On codimension one stability of the soliton for the 1D focusing cubic Klein–Gordon equation. Communications of the AMS, 4(7):230–356, February 2024. ISSN 2692-3688.

Galashin:2024:PCN

[31] Pavel Galashin and Thomas Lam. Positroid Catalan numbers. Communications of the AMS, 4(8):357–386, April 2024. ISSN 2692-3688.

Ganguly:2024:SCD

[32] Shirshendu Ganguly and Alan Hammond. Stability and chaos in dynamical last passage percolation. Communications of the AMS, 4(9):387–479, June 2024. ISSN 2692-3688.

Kra:2024:PES

[33] Bryna Kra, Joel Moreira, Florian Richter, and Donald Robertson. A proof of Erdős’s B + B + t conjecture. Communications of the AMS, 4(10):480–494, August 2024. ISSN 2692-3688.

Lyubich:2024:ACM

[34] Mikhail Lyubich, Jacob Mazor, and Sabyasachi Mukherjee. Antiholomorphic correspondences and mating I: Realization theorems. Communications of the AMS, 4(11):495–547, August 2024. ISSN 2692-3688.

Wright:2024:SCG

[35] Alex Wright. Spheres in the curve graph and linear connectivity of the Gromov boundary. Communications of the AMS, 4(12):548–577, September 2024. ISSN 2692-3688.

Dundas:2024:HHM

[36] Bjørn Dundas, Michael Hill, Kyle Ormsby, and Paul Østvær. Hochschild homology of mod-p motivic cohomology over algebraically closed fields. Communications of the AMS, 4(13):578–606, September 2024. ISSN 2692-3688.

Brenner:2024:NLS

[37] Susanne Brenner, Li yeng Sung, Zhiyu Tan, and Hongchao Zhang. A nonlinear least-squares convexity enforcing C^{0} interior penalty method for the Monge–Ampère equation on strictly convex smooth planar domains. Communications of the AMS, 4(14):607–640, October 2024. ISSN 2692-3688.

Faure:2024:MLA

[38] Frédéric Faure and Masato Tsujii. Microlocal analysis of contact Anosov flows and band structure of the Ruelle spectrum. Communications of the AMS, 4(15):641– 745, November 2024. ISSN 2692-3688.

Chevyrev:2024:SLB

[39] Ilya Chevyrev, Alexey Korepanov, and Ian Melbourne. Superdiffusive limits beyond the Marcus regime for deterministic fast-slow systems. Communications of the AMS, 4(16):746–786, November 2024. ISSN 2692-3688.

Lu:2024:SBG

[40] Jingcheng Lu, Eitan Tadmor, and Anil Zengino lu. Swarm-based gradient descent method for non-convex optimization. Communications of the AMS, 4(17): 787–822, December 2024. ISSN 2692-3688.

Kim:2025:EDS

[41] Dongryul Kim, Hee Oh, and Yahui Wang. Ergodic dichotomy for subspace flows in higher rank. Communications of the AMS, 5(1):1–47, February 2025. ISSN 2692-3688.

Christiansen:2025:LER

[42] T. Christiansen and K. Datchev. Low energy resolvent expansions in dimension two. Communications of the AMS, 5(2): 48–80, April 2025. ISSN 2692-3688.

Gorodetski:2025:NSA

[43] Anton Gorodetski and Victor Kleptsyn. Non-stationary Anderson localization. Communications of the AMS, 5(3): 81–143, April 2025. ISSN 2692-3688.

Kim:2025:FOR

[44] Sang hyun Kim, Thomas Koberda, and J. de la Nuez González. First order rigidity of homeomorphism groups of manifolds. Communications of the AMS, 5(4): 144–194, May 2025. ISSN 2692-3688.

Iyengar:2025:NEU

[45] Srikanth Iyengar, Linquan Ma, Mark Walker, and Ziquan Zhuang. Nonexistence of Ulrich modules over Cohen– Macaulay local rings. Communications of the AMS, 5(5):195–208, May 2025. ISSN 2692-3688.

Hairer:2025:SGP

[46] Martin Hairer and Tommaso Rosati.

Spectral gap for projective processes of linear SPDEs. Communications of the AMS, 5(3):209–283, June 2025. ISSN 2692-3688.

Killip:2025:SCW

[47] Rowan Killip, Thierry Laurens, and Monica Vi an. Scaling-critical well-posedness for continuum Calogero–Moser models on the line. Communications of the AMS, 5 (7):284–320, June 2025. ISSN 2692-3688.

Filippas:2025:UCS

[48] Spyridon Filippas, Camille Laurent, and Matthieu Léautaud. Unique continuation for Schrödinger operators with partially Gevrey coefficients. Communications of the AMS, 5(8):321–391, July 2025. ISSN 2692-3688.

Kazhdan:2025:UES

[49] David Kazhdan and Andrei Okounkov.

On the unramified Eisenstein spectrum. Communications of the AMS, 5(??):392– 456, August 2025. ISSN 2692-3688.